OpenStudy (anonymous):

Find the equation of a line with a given slope and y-intercept : slope-3/5,y-intercept(0,2)

OpenStudy (texaschic101):

y = mx + b m is your slope and b is your y intercept can you solve it now ?

OpenStudy (texaschic101):

just sub in -3/5 for m and 2 for b and what do you get ?

OpenStudy (anonymous):

this is what I got -3x+0

OpenStudy (texaschic101):

no....y = -3/5x + 2

OpenStudy (anonymous):

Oh I know why becos 5y-10=-3x+0 3x+5y=10 so you simplify it

OpenStudy (texaschic101):

I just subbed in the values your slope -3/5 goes in the m position your y intercept, 2, goes in your b position y = mx + b if I was to say you had a slope of 3 and a y intercept of (0,6) then your equation would be : y = 3x + 6 does that make sense ?

OpenStudy (anonymous):

Oh okay

OpenStudy (texaschic101):

or... y - y1 = m(x - x1) slope(m) = -3/5 (0,2) x1 = 0 and y1 = 2 now lets sub y - 2 = -3/5(x - 0) y - 2 = -3/5x - 0 y = -3/5x + 2 standard form would be : Ax + By = C 3/5x + y = 2 3x + 5y = 10 you can use either formula and you will get the same answer

OpenStudy (anonymous):

I just use the same formula that you gave me to solved this : Find the slope intercept form of the equation of a line with the given slope that passes through the given point. slope -1,through (-6,-3) And this is what I got Y=-1 x-9

OpenStudy (texaschic101):

for this one, you need to use y - y1 = m(x - x1) slope(m) = -1 (-6,-3) x1 = -6 and y1 = -3 now lets sub y - (-3) = -1(x - (-6) y + 3 = -1(x + 6) y + 3 = -x - 6 y = -x - 6 - 3 y = -x - 9 (this is the same as y = -1x - 9) you got it :) you are correct

OpenStudy (anonymous):

Thats makes me happy ,I am doing my best to know Algebra ,I am good in basic math and Very good in English ,Economics and business management

OpenStudy (texaschic101):

sounds like you got your hands full. You seem smart, I believe you can handle it. Good job and keep up the good work :)

OpenStudy (anonymous):

Thank you so much and the reason is because it been 7 years that i did math in High and when I was in college in my country I didn't do any math but statistic and economics